# ARBoids: Adaptive Residual Reinforcement Learning With Boids Model for Cooperative Multi-USV Target Defense

### v2

@@ III-E Network Structure
The proposed neural network structure consists of three main modules: state embedding module, decision module, and adapter module. As illustrated in Fig. 2, the state components $s_{i,AT}$ and $s_{i,\text{Boids}}$ are processed independently through single fully-connected (Fc) layers, each with a specific number of hidden units. Meanwhile, $s_{i,D}$ is encoded using a mean observation embedding strategy, following the approach outlined in[28]. Specifically, each observation related to a teammate is passed through an Fc layer, and the resulting features are averaged to produce the final embedding. The Leaky ReLU[29] is selected as the activation function. The three resulting feature vectors are then concatenated and input into the subsequent layers of the network. The critic network employs a similar architecture but includes an additional action embedding layer to process the action input.
@@ III-F Reward Function
Figure 4: Diagram of the components included in the formation reward $r_{\text{form}}$. Here, $\vec{d}_{iA}$ denotes the unit vector from defender $i$ to the attacker, $\vec{d}_{TA}$ is the unit vector from the target to the attacker, and $\vec{d}_{DA}=\sum_{i=1}^{n}\vec{d}_{iA}$ is the cumulative direction vector of all defenders.
@@
where $r_{\text{main}}$ denotes the main task-related reward, $r_{\text{form}}$ encourages formation maintenance, and $r_{\text{collision}}$ penalizes inter-agent collisions.
@@
Specifically, a penalty of $-100$ is assigned if the attacker reaches the target region, while a reward of $100$ is granted upon successful capture by the agent. A helper reward of $50$ is provided if a teammate captures the attacker and the agent is within $3\cdot r_{\text{cap}}$ of the capture location. This design promotes collaboration among defenders.
The formation reward $r_{\text{form}}$ encourages defenders to maintain a symmetric and dispersed defensive formation around the attacker. As illustrated in Fig. 4, $\vec{d}_{iA}$ represents the unit vector pointing from defender $i$ toward the attacker, and $\vec{d}_{DA}=\sum_{i=1}^{n}\vec{d}_{iA}$ represents the cumulative direction vector of all defenders. The reward is computed as
@@
The collision penalty $r_{\text{collision}}$ discourages interference between defenders and is defined as
$r_{\text{collision}}=\begin{cases}-50,&\text{if}\ d_{i,j}\leq r_{\text{collision}},\\ 0,&\text{otherwise}.\end{cases}$ (18)
No additional inter-agent attraction or spacing constraints are imposed, as such mechanisms may hinder the learning of an effective defensive strategy. Overall, this reward structure is designed to balance individual performance, collaborative behavior, and formation integrity, ultimately enhancing the overall robustness of the learned policy.
@@ III-G Curriculum Learning
Curriculum learning [30] is employed to enhance training stability and mitigate the risk of converging to suboptimal local minima. In this work, we progressively increase the TDP task difficulty to facilitate effective exploration during the early stages of training. The primary factor influencing task difficulty is the attacker’s agility level, $\mathcal{L}_{\text{agi}}$. At the beginning of each training episode, $\mathcal{L}_{\text{agi}}$ is randomly sampled from a uniform distribution as follows,
@@
where ¯$\bar{\mathcal{L}}_{\text{agi}}$ represents the average agility level. To initiate learning with manageable complexity, ¯$\bar{\mathcal{L}}_{\text{agi}}$ is initially set to a relatively low value (¯$\bar{\mathcal{L}}_{\text{agi}}=2.0$), enabling agents to effectively learn the fundamental dynamics of the task. Over the course of training, ¯$\bar{\mathcal{L}}_{\text{agi}}$ is increased by $0.25$ every $2.5\times 10^{5}$ training steps, which corresponds to one-quarter of the total training duration. This progressive increase in task difficulty encourages the agents to develop robust and generalizable defensive strategies, rather than overfitting to simplified scenarios with limited attacker agility.
IV Experiments and Discussions
@@ IV-A1 Experiment Parameters
[…]
@@ IV-A2 Attacker’s Policy
The attacker employs an improved artificial potential field (APF) strategy[32], which combines both attractive and repulsive forces. The target region generates an attractive force that pulls the attacker toward the target, while each defender within the attacker’s sensing range induces a repulsive force to steer it away. Due to frequent collinearity between the target and the defenders, the attacker is prone to being trapped in local extrema. To mitigate this issue, a rotational repulsive force is incorporated, as described in [32], enabling the attacker circumvent defenders and escape such disadvantageous configurations effectively.
IV-A3 Gazebo Environment
[…]
@@ IV-B Experimental Results
Figure 5: Snapshots (left) and trajectories (right) from a representative Gazebo VRX experiment deploying the proposed ARBoids method. VRX is a high-fidelity marine simulator developed by Open Robotics [34], and the USV model follows the full dynamics in (1). The attacker, with an agility level set to $\mathcal{L}_{\text{agi}}=2.25$, attempts to penetrate the dock area. The timestamp $t$ is displayed in the upper-left corner of each snapshot. Red and blue arrows denote the heading directions of the attacker and defender, respectively.
Figure 6: Comparison of learning curves for ARBoids, RP, and SAC. Success rates (SRs) are recorded every $5000$ steps to evaluate the learning performance of each method. The shaded areas denote the standard deviation across 5 independent runs, reflecting the stability of the results. The proposed ARBoids exhibits rapid convergence to a high-performing policy, substantially outperforming both SAC and RP in terms of final SR and learning efficiency.
@@ IV-C1 Benchmark Methods
To demonstrate the advantages of our proposed method, we quantitatively evaluate its performance by comparing its success rate (SR) and convergence speed against three benchmark approaches: the Boids model, the standard residual policy (RP), and a vanilla DRL policy trained using the SAC algorithm. For the Boids model, all weights are configured identically to those used in our method to ensure a fair comparison. In the standard residual policy, action commands are computed as described in (4). For the vanilla SAC policy, the Boids-related state component $s_{i,\text{Boids}}$ is excluded from the state space. Each method is evaluated for 100 trials under identical experimental conditions.
@@ IV-C2 Comparison Results
We track the SRs of three learning-based methods throughout the training process. Each method is evaluated every $5000$ training steps in a scenario where the attacker’s agility level is fixed at $\mathcal{L}_{\text{agi}}=2.0$. As shown in Fig. 6, the vanilla SAC algorithm fails to converge to an effective policy, with its SR stagnating around 20% even after a million steps, exhibiting minimal improvement over time. In contrast, both the standard RP and ARBoids methods benefit from the Boids component, which provides a strong initial performance of approximately 45% and facilitates exploration through increasing the frequency of positive rewards. However, RP shows limited improvement as training progresses, with its SR plateauing around 55% after 0.1 million steps. This suggests a restricted capacity for further optimization, likely due to the absence of adaptive mechanisms. The proposed ARBoids method, by comparison, demonstrates rapid convergence to an effective strategy, achieving an SR of 90% within just 0.35 million steps. This represents a significant 35% absolute improvement over RP and 70% over SAC. These results highlight the superior sample efficiency, training stability, and final performance achieved by ARBoids, validating the effectiveness of its adaptive hybrid architecture.
IV-C3 Robustness to Attacker Agility Levels
@@
We evaluate the SRs of all comparison methods under varying attacker agility levels, with $n=3$ defenders. As illustrated in Fig. 7(a), the SRs of all methods decline as the attacker’s agility increases, reflecting the growing difficulty of maintaining effective defense under more agile adversarial behavior. Across all agility levels, the proposed ARBoids consistently and significantly outperforms the comparison methods, with the performance gap widening as the problem becomes more challenging. Specifically, while the Boids model and the standard RP achieve SRs close to 80% at $\mathcal{L}_{\text{agi}}=1.5$, their performance deteriorates sharply to below 40% when $\mathcal{L}_{\text{agi}}\geq 2.5$. In contrast, ARBoids exhibits a relatively modest performance drop of approximately 18%, maintaining high SRs even under challenging conditions. These findings underscore the robustness and generalizability of ARBoids across a wide range of attacker agility levels—an essential capability for real-world deployment scenarios where the attacker’s agility is unknown or variable.
IV-C4 Scalability With Respect to Team Size
@@
We further evaluate the scalability of the proposed ARBoids method under varying numbers of defenders. Although the policy is exclusively trained in a 3-defender setting, it exhibits inherent adaptability to different team sizes ($n$) due to the utilization of the mean observation embedding technique [28]. The corresponding SRs are illustrated in Fig. 7(b). As the number of defenders increases, the defense scenario becomes more complex. While additional defenders expand coverage and interception capability, they also introduce challenges such as increased collision risk and higher coordination demands. The proposed ARBoids maintains high SRs for $n=3$ and $n=4$, demonstrating robustness in typical deployment scenarios. However, performance begins to degrade for $n\geq 5$, indicating that densely coordinated environments may require enhanced collision avoidance strategies to sustain optimal performance.
@@ V Conclusion
This letter investigated the target defense problem involving an agile attacking USV and a team of homogeneous defending USVs. We proposed a novel defending control strategy, ARBoids, which integrated the SAC algorithm with the biologically inspired Boids model. The Boids model served as a baseline coordination policy, while an SAC agent learned a residual policy to optimize defenders’ actions. An adapter module dynamically balanced these contributions. A tailored reward function and curriculum learning enhanced training stability and generalization. Extensive experiments conducted in a high-fidelity marine simulation environment demonstrated that ARBoids significantly outperformed selected benchmark methods. Future work will enhance the ARBoids framework by incorporating advanced attacker policies, realistic perception models, and environmental factors like obstacles and waves. We also plan to develop collision avoidance techniques and validate the approach in real-world experiments.

### v3

@@ III-E Network Structure
The proposed neural network structure consists of three main modules: state embedding module, decision module, and adapter module. As illustrated in Fig. 2, the state components $s_{i,AT}$ and $s_{i,\text{Boids}}$ are processed independently through single fully-connected (Fc) layers, each with a specific number of hidden units. Meanwhile, $s_{i,D}$ is encoded using a mean observation embedding strategy, following the approach outlined in[huttenrauch2019deep]. Specifically, each observation related to a teammate is passed through an Fc layer, and the resulting features are averaged to produce the final embedding. The Leaky ReLU is selected as the activation function. The three resulting feature vectors are then concatenated and input into the subsequent layers of the network. The critic network employs a similar architecture but includes an additional action embedding layer to process the action input.
@@ III-F Reward Function
Specifically, a penalty of 100-100 is assigned if the attacker reaches the target region, while a reward of $100$ is granted upon successful capture by the agent. A helper reward of $50$ is provided if a teammate captures the attacker and the agent is within $3\cdot r_{\text{cap}}$ of the capture location. This design promotes collaboration among defenders.
The formation reward $r_{\text{form}}$ encourages defenders to maintain an effective defensive formation around the attacker. As illustrated in Fig. 3(b), $\vec{d}_{iA}$ represents the unit vector pointing from defender $i$ toward the attacker, and $\vec{d}_{DA}=\sum_{i=1}^{n}\vec{d}_{iA}$ represents the cumulative direction vector of all defenders. The formation reward is computed as
@@
To discourage interference among defenders, the collision penalty is defined as $r_{\text{collision}}=-50$ if a collision occurs ($d_{i,j}\leq r_{\text{collision}}$), and $r_{\text{collision}}=0$ otherwise. No additional inter-agent attraction or spacing rewards are applied, as such mechanisms are inherently similar to the Boids virtual forces, leading to redundancy. Besides, the reward components are carefully scaled to reflect their relative importance. Specifically, $r_{\text{main}}$ and $r_{\text{collision}}$ are prioritized to dominate the optimization process, while $r_{\text{form}}$ serves as auxiliary guidance.
@@ III-G Curriculum Learning
Curriculum learning [bengio2009curriculum] is employed to enhance training stability and mitigate the risk of converging to suboptimal local minima. In this work, we progressively increase the TDP task difficulty to facilitate effective exploration during the early stages of training. The primary factor influencing task difficulty is the attacker’s agility level, $\mathcal{L}_{\text{agi}}$. At the beginning of each training episode, $\mathcal{L}_{\text{agi}}$ is randomly sampled from a uniform distribution as follows,
@@
where $\bar{\mathcal{L}}_{\text{agi}}$ represents the average agility level. To initiate learning with manageable complexity, $\bar{\mathcal{L}}_{\text{agi}}$ is initially set to a relatively low value ($\bar{\mathcal{L}}_{\text{agi}}=2.0$), enabling agents to effectively learn the fundamental dynamics of the task. Over the course of training, $\bar{\mathcal{L}}_{\text{agi}}$ is increased by $0.25$ every $2.5\times 10^{5}$ training steps, which corresponds to one-quarter of the total training duration. This progressive increase in task difficulty encourages the agents to develop robust and generalizable defensive strategies, rather than overfitting to simplified scenarios with limited attacker agility.
IV Experiments and Analysis
@@ IV-A1 Experiment Parameters
The simulation parameters are configured as follows. The capture radius $r_{\text{cap}}$ is set to $5\,\text{m}$. The target radius $r_{0}$, the target’s sensing range $\rho_{T}$, and the attacker’s sensing range $\rho_{A}$ are set to $15\,\text{m}$, $60\,\text{m}$, and $15\,\text{m}$, respectively. The maximum forward thrust of each defending USV $\tau^{\text{max}}_{D}$ is set to $1000\,\text{N}$, while the maximum reverse thrust $\tau^{\text{min}}_{D}$ is set to $500\,\text{N}$. The total simulation time is $T_{\text{total}}=60\,\text{s}$, and the action interval is $T_{\text{action}}=0.2\,\text{s}$. The Boids model weights are configured with separation $k_{\text{sep}}=10$, alignment $k_{\text{ali}}=0.1$, cohesion $k_{\text{coh}}=0.1$, and attraction $k_{\text{att}}=0.5$. During training, the number of defenders is fixed at $3$. The SAC algorithm is adopted with a learning rate of $1\times 10^{-4}$, a batch size of $4096$, a total of $1\times 10^{6}$ training steps, and a discount factor of $0.99$.
@@ IV-A2 Attacker’s Policy
The attacker employs an improved artificial potential field (APF) strategy[zhang2024integrated], which combines both attractive and repulsive forces. The target region generates an attractive force that pulls the attacker toward the target, while each defender within the attacker’s sensing range induces a repulsive force to steer it away. Furthermore, we investigate the robustness of our approach against learning-based attacker policies. An alternating learning framework is adopted, in which the attacker and defender alternately learn their respective policies. During each phase, one side’s policy is fixed while the other undergoes training, enabling iterative mutual adaptation and performance improvement. The attacker’s policy is optimized using the SAC algorithm, guided by a deliberately designed reward function that balances the objectives of reaching the target and avoiding defenders.
IV-A3 Training Environment
To enable large-scale DRL rollouts with stable wall-clock time, the training environment employs a custom numerical simulator based on the planar 3-DoF dynamic model. This model simplifies the vessel’s dynamics (1) to the horizontal plane, focusing on surge, sway, and yaw rate components, which retains the key maneuvering dynamics necessary for the tasks. The environment and models are implemented in Python, with PyTorch used for neural network development. Training and evaluations are performed on a workstation equipped with an Intel i7-14650HX CPU and an NVIDIA GeForce RTX 4070Ti GPU. Under this setup, running 1 million training steps with our method requires approximately 7.6 hours of wall-clock time.
@@ IV-B Experimental Results
Figure 4: Snapshots (left) and trajectories (right) from a representative Gazebo VRX experiment deploying the proposed ARBoids method. VRX is a high-fidelity marine simulator and the USV model follows the dynamics in (1). The attacker, with an agility level of $\mathcal{L}_{\text{agi}}=2.25$, tries to penetrate the dock area. The timestamp $t$ is displayed in the upper-left corner of each snapshot. Red and blue arrows denote the heading directions of the attacker and defender, respectively.
Figure 5: Learning curves comparing our full method (ARBoids), ablations (w/o formation reward, FR; w/o curriculum learning, CL), and baselines (RP, SAC, Boids). Success rates are measured every 5,000 steps to assess learning progress. Shaded regions indicate the standard deviation across 5 independent runs. ARBoids substantially outperforms all baselines, and removing either FR or CL results in a pronounced degradation in performance.
@@ IV-C1 Benchmark Methods
To demonstrate the advantages of our proposed method, we quantitatively evaluate its performance by comparing its success rate (SR) and convergence speed against three benchmark approaches: the Boids model, the standard RP, and a vanilla DRL policy trained using the SAC algorithm. For the Boids model, all weights are configured identically to those used in our method to ensure a fair comparison. In the standard residual policy, action commands are computed as described in (4). For the vanilla SAC policy, the Boids-related state component $s_{i,\text{Boids}}$ is excluded from the state space. Each method is evaluated for 100 trials under identical experimental conditions.
@@ IV-C2 Comparison Results
We track the SRs of three learning-based methods—SAC, RP and ARBoids along with two ablations of ARBoids without the formation reward (w/o FR) and without curriculum learning (w/o CL). Each method is assessed every 5,000 steps under a fixed attacker agility level of $\mathcal{L}_{\text{agi}}=2.0$. As shown in Fig. 5, vanilla SAC fails to converge, with SR stagnating near 20% even after 1 million steps. Both RP and ARBoids benefit from the Boids prior, yielding a strong initial SR of about 45% by promoting exploration via more frequent positive rewards. However, RP quickly plateaus around 55% after 0.1 million steps, indicating limited capacity for further improvement. By contrast, ARBoids converges rapidly and reliably, reaching 90% SR within 0.35 million steps—an absolute gain of 35% over RP and 70% over SAC. The ablations confirm the contributions of both components. Removing either FR or CL leads to a significant performance degradation. Overall, ARBoids achieves superior sample efficiency, stability, and final SR, validating the effectiveness of its adaptive hybrid design.
IV-D Robustness Against Attacker Agility Levels
@@
We evaluate the SRs of our method under varying attacker agility levels, with $n=3$ defenders. As illustrated in Fig. 6(a), the SRs of all methods decline as the attacker’s agility increases, reflecting the growing difficulty of maintaining effective defense under more agile adversarial behavior. Across all agility levels, the proposed ARBoids consistently and significantly outperforms the comparison methods, with the performance gap widening as the problem becomes more challenging. Specifically, while the Boids model and the standard RP achieve SRs close to 80% at $\mathcal{L}_{\text{agi}}=1.5$, their performance deteriorates sharply to below 40% when $\mathcal{L}_{\text{agi}}\geq 2.5$. In contrast, ARBoids exhibits a relatively modest performance decrease of approximately 18%, maintaining high SRs even under challenging conditions. These findings underscore the robustness and generalizability of ARBoids across a wide range of attacker agility levels—an essential capability for real-world deployment scenarios where the attacker’s agility is unknown or variable.
IV-E Generalization to Unseen Team Size
@@
We evaluate the generalization capability of ARBoids under unseen defender team sizes. Although the policy is exclusively trained with three defenders, it successfully generalizes across different team sizes ($n$), enabled by the mean observation embedding technique [huttenrauch2019deep]. The corresponding SRs are illustrated in Fig. 6(b). As the number of defenders increases, the defense scenario becomes more complex. While additional defenders expand coverage and interception capability, they also introduce challenges such as increased collision risk and higher coordination demands. ARBoids maintains high SRs for $n=3$ and $n=4$, demonstrating robustness in typical deployment scenarios. However, performance begins to degrade when $n>4$, indicating that densely coordinated environments require stronger coordination and collision avoidance mechanisms, which the current policy, trained only with three defenders, cannot fully provide. To address this, we plan to explicitly train ARBoids on larger defender teams and enhance its scalability through techniques such as value decomposition networks and behavior prediction for collision avoidance.
IV-F Robustness Against Learning-Based Attacker
Figure 7: Learning curves of the defender and attacker during alternating training. The curves illustrate the defender success rate (Defender SR), collision rate, and attacker success rate (Attacker SR) as functions of training steps. The blue-shaded regions indicate the defender’s learning phases, whereas the red-shaded regions correspond to the attacker’s learning phases.
We further evaluate the robustness of ARBoids against a learning-based attacker by training the defender and attacker alternately over five phases, totaling 2.5 million training steps. The learning curves of both agents during this process are shown in Fig. 7. During each attacker learning phase, the defender’s SR temporarily decreases as the attacker learns to exploit weaknesses in the defender’s policy. However, it quickly recovers and improves in the subsequent defender learning phase, demonstrating the effectiveness of our method in adapting to evolving attacker strategies through fine-tuning. Additionally, the defenders’ collision rate starts relatively high during the early stages of training but steadily decreases as training progresses. This indicates that the defenders not only develop policy-specific counter-strategies but also acquire general collision-avoidance capabilities, highlighting their improved adaptability. However, we observe that the final defender policy derived from this framework exhibits a slightly reduced SR when tested against the original APF-based attacker, revealing that defender policies still have a tendency to overfit to specific attack strategies encountered during training.
@@ V Conclusion
This letter investigated the target defense problem involving an agile attacking USV and a team of homogeneous defending USVs. We proposed a novel defending control strategy, ARBoids, which integrated the SAC algorithm with the biologically inspired Boids model. The Boids model served as a baseline coordination policy, while an SAC agent learned a residual policy to optimize defenders’ actions. An adapter module dynamically balanced these contributions. A tailored reward function and curriculum learning enhanced training stability and generalization. Extensive experiments conducted in a high-fidelity marine simulation environment demonstrated that ARBoids significantly outperformed selected benchmark methods. Future work will enhance the generalization of ARBoids across diverse and unseen attack strategies by further exploring the alternating training with techniques such as adversarial policy ensembles. We also plan to develop advanced collision avoidance techniques, investigate scaling limits, and validate our approach in real-world experiments.